Circular No. 6272

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GALACTIC CENTER G. J. Fishman, Marshall Space Flight Center (MSFC), NASA; C.

Kouveliotou, Universities Space Research Association and MSFC; J. van Paradijs, University of Alabama at Huntsville (UAH) and University of Amsterdam; B. A. Harmon, MSFC; W. S. Paciesas and M. S. Briggs, UAH; and J. Kommers and W. H. G. Lewin, Massachusetts Institute of Technology, report observations from the Compton Observatory’s BATSE experiment: “A transient x-ray source is being observed that emits bursts at energies between 10 and 50 keV from the general direction of the Galactic Center. Over 200 bursts with typical widths between 8 and > 30 s have been seen from the source since Dec. 2. The rate of the x-ray bursts has declined from a maximum of ∼ 18 events per hour on Dec. 2.7 UT to a current average rate of ∼ 1.2 events per hour (corrected for live-time and earth occultation). The average burst peak flux is ∼ 5 × 10^{−8} erg cm^{−2} s^{−1} and the typical fluence is ∼ 5 × 10^{−7} erg/cm^{2} per burst in the energy band 20–40 keV. No steady emission is seen from the source. The spectra of the events are similar to bursts from soft γ-ray repeaters, but much harder than those of type-II bursts from the rapid burster 1730–335. The most probable location of the source is centered at α = 17^{h}52^{m}, δ = −23^{o}, with an error radius of ∼ 6^{o} (1σ). The behavior reported here is unlike that previously seen from any x-ray or γ-ray transient source.”

4U 0115+634 S. Sazonov and R. Sunyaev, Space Research Institute, Moscow, on

behalf of the Granat/WATCH team, report: “The transient x-ray pulsar 4U 0115+634 is undergoing a strong outburst. The flux from 8 to 20 keV increased at a constant rate of ∼ 100 mCrab/day from 60 ± 50 mCrab on Nov. 18 to 670 ± 60 mCrab on Nov. 25; it remained at that level in subsequent WATCH observations, on Nov. 26–29 and Dec. 5–8. This outburst already reported by the BATSE team (IAUC 6266) is the second major outburst observed from the source with WATCH during the period 1989–1995. It is similar in temporal behavior and intensity to the previous one observed in 1990 Feb.–Mar. by Granat/WATCH and Ginga. These two outbursts are separated by 5.8 yr, which is consistent with the 3-yr quasi-periodicity previously known to exist in 4U 0115+634 (Tamura et al. 1992, Ap.J. 389, 676).”

1995 December 12 Daniel W. E. Green